OpenStudy (anonymous):

Find the range of the third side of a triangle given the other 2 sides are 6 and 12.

OpenStudy (anonymous):

\[x²=6^{2}+12^{2}-2.6.12.\cos(\theta)\] \[-1\le \cos(\theta) \le 1\]

OpenStudy (anonymous):

answer choices not enough information 12 < x < 24 36 < x < 144 6 < x < 18 3 > x < 6